Find the equation of the line with slope 6 that goes through the point (6,4). Write your answer in slope-intercept form.
Answer by checkley77(12844)

Y=mX+b
4=6*6+b
4=36+b
b=4-36
b=-32 ANSWER FOR THE Y INTERCEPT.
Y=6X-32 LINE EQUATION ANSWER.